OpenStudy (anonymous):

Find the slope of a line parallel to the line with equation 3x – 2y = 12.

OpenStudy (anonymous):

3x - 2y = 12 -2y = 12 - 3x 2y = 3x - 12 y = 2x/3 - 12 Parallel lines: m1 = m2 that means the slope of the parallel line is equal to this one. which is 2/3

OpenStudy (anonymous):

\[3x - 2y = 12\] \[2y = 3x - 12\] \[y = \frac{ 3 }{ 2 }x - \frac{ 12 }{ 2 }\] on comparing it with \[y = mx + c\] we get \[m = \frac{ 3 }{ 2 }\] when the lines are parallel the slopes are equal therefore m = 3/2

OpenStudy (moonlitfate):

First put the equation in y = mx + b form; M is equal to the slope of the line. \[y = \frac{ -3x }{ 2 } + 4\] The slope is -3x/2. A line that's parallel has the same slope.

OpenStudy (moonlitfate):

*3/2

OpenStudy (anonymous):

@MoonlitFate dont you think you are wrong ? b = 6 i guess and again you are wrong in sign conventions ;)

OpenStudy (anonymous):

okay thanks so much i understand can you help me with this one

OpenStudy (anonymous):

Find the slope of a line perpendicular to the line with equation 4x + 2y = -8

OpenStudy (anonymous):

ya sure but where is my medal :D

OpenStudy (moonlitfate):

@miteshchvm Oh, you're right. Haha, guess midnight is a bit too late to be doing math. Careless mistakes. :p

OpenStudy (anonymous):

\[4x +2y = -8\] \[2y = -4x - 8\] \[y = -2x -4\] on comparing we get m = -2 here the slopes are perpendicular hence let the slope of the other line be 'n' condition of perpendicular lines slopes is \[m \times n = \left( -1 \right)\] hence \[-2 \times n = -1\] \[n = \left( \frac{ -1 }{ -2 } \right)\] hence \[n = \left( \frac{ 1 }{ 2 } \right)\]

OpenStudy (anonymous):

thanks